# Learning General Continuous Constraint from Demonstrations via Positive-Unlabeled Learning

### v1

@@ Abstract
Planning for a wide range of real-world tasks necessitates to know and write all constraints. However, instances exist where these constraints are either unknown or challenging to specify accurately. A possible solution is to infer the unknown constraints from expert demonstration. The majority of prior works limit themselves to learning simple linear constraints, or require strong knowledge of the true constraint parameterization or environmental model. To mitigate these problems, this paper presents a positive-unlabeled (PU) learning approach to infer a continuous, arbitrary and possibly nonlinear, constraint from demonstration. From a PU learning view, We treat all data in demonstrations as positive (feasible) data, and learn a (sub)-optimal policy to generate high-reward-winning but potentially infeasible trajectories, which serve as unlabeled data containing both feasible and infeasible states. Under an assumption on data distribution, a feasible-infeasible classifier (i.e., constraint model) is learned from the two datasets through a postprocessing PU learning technique. The entire method employs an iterative framework alternating between updating the policy, which generates and selects higher-reward policies, and updating the constraint model. Additionally, a memory buffer is introduced to record and reuse samples from previous iterations to prevent forgetting. The effectiveness of the proposed method is validated in two Mujoco environments, successfully inferring continuous nonlinear constraints and outperforming a baseline method in terms of constraint accuracy and policy safety.
Keywords: Constraints inference, Inverse reinforcement learning, Learning from expert demonstrations
1 Introduction
@@
Planning for many robotics and automation tasks also requires knowing constraints explicitly, which define what states or trajectories are allowed or must be avoided [1, 2]. Sometimes these constraints are initially unknown or hard to specify mathematically, especially when they are general, continuous, or inherent to an expert’s preference and experience. For example, human drivers may determine an implicit minimum distance from other cars based on traffic conditions, traffic rules, and even weather. To learn a driving policy matching human behaviors, an explicit constraint should be inferred somehow, e.g., from existing human demonstration sets.
A typical approach to recover the underlying constraint is through Inverse Constrained Reinforcement Learning (ICRL) [3, 4, 5]. It originates from Inverse Reinforcement Learning (IRL), a method to infer a reward function by observing the demonstrations of an expert [6]. In contrast to IRL, ICRL is developed exclusively to infer constraints instead of reward functions. A current major challenge of ICRL is that most methods are limited to recovering a linear constraint function or a nonlinear constraint with known parameterization. In this study, we will focus on this challenge and aim at learning a continuous (possibly nonlinear) constraint function on continuous state-action spaces. To see this, next we will review the existing constraint inferring methods, summarize their limitations, and explain the motivation of this work.
Figure 1: The framework of the proposed method. It alternates between two steps: learning policy and learning constraint. The first step updates the policy network by maximizing the reward subject to the already learned constraint. In second step, we utilize an PU learning approach to infer the unknown infeasible areas from both the demonstrated trajectories and the sampled potentially unsafe trajectories.
@@
Related Works: Constraint inference through ICRL has drawn more and more attention since 2018. Chou et al. [7] first explores how to infer constraints given an environment model and a demonstration set. They assume that all possible trajectories that could earn higher rewards than the demonstration must be constrained in some way, or the expert could have passed those trajectories. In practice, they use hit-and-run sampling to obtain such higher reward trajectories and solve an integer program to recover the constrained states in a tabular environment. Scobee and Sastry [8] formalizes this idea by casting the problem in the maximum likelihood inference framework, which is also prevalent in the IRL domain. They introduce a Boltzmann policy model, where the likelihood of any feasible trajectory is assumed to be proportional to the exponential return of the trajectory, while the likelihood of any infeasible trajectory is 0. Then a greedy algorithm is proposed to add the smallest number of constraints that maximize the likelihood of the demonstrations. This framework has the advantage of being able to work with sub-optimal demonstrations. Alternatively, Vazquez-Chanlatte et al. [9] formulate a maximum a posterior probability inference problem to learn non-Markovian constraint, or task specifications. Further, Glazier et al. [10] and McPherson et al. [11] extend the maximum entropy framework from a deterministic setting into a stochastic setting, with soft or probabilistic constraints. In [10], they learn a residual reward (i.e., penalty) from demonstration using maximum entropy IRL. Assuming the penalty value follows a logistic distribution, the probability of constraint over features can be recovered. In [11], a causal entropy likelihood function is formulated and optimized via a modified version of soft Bellman backup, which is challenging to scale to continuous state-action spaces.
@@
Unfortunately, all the methods mentioned above only apply to systems with discrete finite state spaces. To overcome this drawback, recent papers have made a few attempts to extend this to continuous cases. Stocking et al. [12] proposes an algorithm that learns a neural network policy via deep RL and generates high return trajectories with the policy network. Then the constraint is again added based on the maximum likelihood principle. However, like the previous methods, this method recovers constraints from a given constraint set. Besides, the method only works with a categorical action policy, not applied to a continuous policy. Recently, Malik et al. [3] proposes Maximum Entropy Constraint Learning (MECL) algorithm. It not only approximates the policy with a network but also learns a constraint network, a continuous function that can represent more general constraints. This constraint network is optimized by making a gradient ascent on the maximum likelihood objective function. Although theoretically the method can arbitrary constraint on continuous state spaces and constraints, in practice it was only applied to recover plane constraints, e.g. $x\geq-3$ [3], and, as we show later in the paper, performs relatively poorly at learning more complex constraints. This is partly due to an unrecognized problem of constraint forgetting that we discuss later in section 2.4. In parallel, in Inverse Optimal Control community, another line of research based on KKT optimality condition has been developed to learn general continuous constraints [13, 14], but these methods require a closed-form model and its derivative, which are not always available in reality.
In summary, despite impressive advances to solve ICRL, it still remains to be shown that ICRL can be used to learn continuous nonlinear constraint functions from a small sample of demonstrations in a model-free manner.
@@
To overcome the aforementioned challenge, we propose a novel positive-unlabeled approach to learn an arbitrary and continuous constraint function in continuous state-action spaces. Our method is inspired by a machine learning subarea Positive-Unlabeled (PU) learning [15], and to best of our knowledge, this work is the first to handle the constraint inference problem from a PU learning perspective. We treat all data in demonstrations as positive (feasible) dataset, and learn a policy to generate many potentially unsafe trajectories, which contains unlabeled states that could be feasible or infeasible. Then a postprocessing PU learning technique is applied to identify the truly infeasible states from the two dataset. It makes an assumption on data distribution, and synthesises the feasible-infeasible classifier (i.e., constraint model) from a feasible-unlabeled classifier with a moved classification threshold.
To generate potentially unsafe trajectories, we also maintain a constrained RL policy that always maximizes the reward function subject to the current constraint function. If the policy achieves a higher reward than the demonstration assumed to be (sub-)optimal, we believe it implies that the policy must have explored certain unknown unsafe states to win such a high reward [7].
@@
Our constraint learning paradigm is phrased as an iterative framework, see Fig. 1. At each iteration, we first train a policy maximizing the reward subject to the constraints. Then, a set of higher-reward trajectories are generated by sampling from the policy, which are further used along with the demonstrations to infer constraints via PU learning technique. In the next iteration, the new policy is updated with respect to the new constraints. To prevent forgetting the previously learned constraints, a memory mechanism is introduced, where some most representative infeasible states are recorded and used for training in the following interactions.
2 Method
2.1 Preliminaries and Problem Statements
@@
For a Markov decision process (MDP) [16, 17], states and actions are denoted by $s\in\mathcal{S}$ and $a\in\mathcal{A}$. $\gamma$ denotes the discount factor, and the real-valued reward is denoted by $r(s,a)$. Note that in ICRL, reward is usually assumed to be given [3, 7]. A trajectory $\tau=\{s_{1},a_{1},\ldots,s_{T},a_{T}\}$ contains a sequence of state-action pairs in one episode. For the trajectory $\tau$, the total discounted reward (return) is defined as $r(\tau)=\sum_{t=1}^{T}\gamma^{t}r(s_{t},a_{t})$. A policy, the mapping between states and actions, is denoted by $\pi(a|s)$. Note that different from [14, 13], we do not make the assumption of possessing a closed-form model or its derivatives of the transition dynamics. We only assume a simulator is available that can simulate the dynamics and return the next state and reward, but no information about the true constraint is included.
@@
The true constraint set $\mathcal{C^{*}}$ consists of all the actually infeasible states $\mathcal{C^{*}}=\{s\in\mathcal{S}|s\texttt{ is actually infeasible}\}$. Note that for simplicity of explanation, we only consider a state constraint, but it can be easily extended to state-action constraints by augmenting the state with the action to form an augmented state. We aim to recover the true constraint set from the demonstration. Suppose we have collected a set of demonstrated trajectories $\mathcal{D}=\{\tau^{i}\}^{M_{d}}_{i=1}$ generated from an expert $\pi^{*}$ navigating in the true environment. We assume that the expert $\pi^{*}$ maximizes the return $J(\pi)=\mathbb{E}_{\tau\sim\pi}\{r(\tau)\}$ while never visiting the truly infeasible states. To represent and learn an arbitrary constraint set in continuous state space, we define a continuous constraint function $\zeta_{\theta}(s)\in(0,1)$ and its induced constraint set $\mathcal{C}_{\theta}=\{s\in\mathcal{S}|\zeta_{\theta}(s)\leq d\}$, where $\theta$ is the function parameter and $d$ is a manually specified threshold. The constraint function $\zeta_{\theta}$ can be regarded as the probability that a state is feasible, while $d$ serves as the decision threshold. We further define a binary constraint indicator function $c_{\theta}(s)$ to represent the feasibility of a state.
$c_{\theta}(s)=\begin{cases}0&\text{if $\zeta_{\theta}(s)>d$ (feasible)}\\ 1&\text{if $\zeta_{\theta}(s)\leq d$ (infeasible)}\\ \end{cases}$ (1)
2.2 Constraint Inference as Positive-Unlabeled Learning
@@
As discussed before, we infer the underlying constraint by contrasting the demonstration with a high-reward-winning policy. In each iteration, we first sample a set of high-reward trajectories $\mathcal{P}=\{\tau_{p}^{i}\}^{M_{p}}_{i=1}$ by performing current policy $\pi_{\phi}$ (discussed later in 2.3). Each sampled trajectory consists of sequential state-action pairs $\tau_{p}=\{s_{j},a_{j}\}_{j=1}^{N_{i}}$. We speculate that the higher-reward trajectory wins high reward by violating some unknown constraints. However, it remains unclear which specific state(s) within the trajectory has violated the constraint. In another word, trajectory $\tau_{p}$ consist of both feasible states and infeasible states but both remains unlabeled. In contrast, it is certain that all states on the demonstrated trajectories are labeled as feasible. Our goal is to classify each single state as feasible or infeasible by learning from a batch of fully labeled feasible samples and another batch of unlabeled samples.
This insight inspires us to formulate constraint inference as a positive-unlabeled learning problem, which learns from only a positive dataset and an unlabeled dataset containing both positive and negative samples [15]. Within our framework, feasible states are designated as positive samples. Consequently, the demonstrations serve as positive samples, while the policy offers unlabeled samples with unknown feasibility. Any datapoint can be represented as a set of triplets $(s,c,l)$ with $s$ the state, $c$ the class (1 for feasible), and $l$ a binary variable representing whether the tuple is labeled. If a sample is labeled (i.e., $l=1$), it must originate from the demonstration and is sure to be feasible, i.e., $Pr(c=1|l=1)=1$. A key quantity from PU learning is the label frequency $f$, which is defined as the proportion of positive samples that are labeled within the entire dataset:
@@
In the constraint learning context, the label frequency $f$ implies the sparsity of the truly feasible states in the state space. A high value $f$ suggests that the majority of truly feasible states are distributed on the demonstrated trajectories and thus labeled, while only a few amount of feasible states are on the higher-reward trajectory and thus unlabeled. Subsequently, we introduce an assumption regarding the labeling mechanism [15]:
@@
Subsequently we introduce a postprocessing approach [18] for uncovering the unknown infeasible regions due to its 1) inherent simplicity and 2) compatibility with various models, including neural networks. Given the SCAR assumption, the probability of an sample being labeled is directly proportional to the probability of that sample being positive:
@@
Consequently, it is evident that a labeled-unlabeled classifier, trained to predict $Pr(l|s)$ by treating unlabeled data as negative and labeled data as positive, can be directly employed to forecast the class probabilities $Pr(c=1|s)$ and predict the class [18]. Alternatively, noticing that $Pr(c=1|s)>0.5$ is equivalent to $Pr(l=1|s)>0.5f$, the labeled-unlabeled classifier can be repurposed as a feasible-infeasible classifier by directly adjusting the decision threshold from $d=0.5$ to $d=0.5f$.
@@
Once trained, this classifier can be directly employed as a feasible-infeasible classifier by setting the decision threshold to $d=0.5f$ as in (7).
@@
The subsequent work involves determining the label frequency $f$. Estimating $f$ is a central concern in PU learning, with diverse methods such as partial matching and kernel embedding being proposed [15]. In this work, we suggest two approaches. A conventional and straightforward approach [19] uses the insight that $Pr(l=1|s)=fPr(c=1|s)$, which is equal to the label frequency $f$ when the true class probability is $Pr(c=1|s)=1$. Suppose there will be state for which $Pr(c=1|s)\approx 1$, the label frequency can be hence estimated as $f=\max_{s\in\mathcal{D}}\zeta_{\theta}(s)$. Alternatively, recall that $f$ indicates the density of infeasible states in the state space, it can be reviewed as a hyper-parameter representing the user’s belief over the constraint density. A higher value of $f$ suggests a denser distribution of infeasible states. In our following experiments, we adopt the latter perspective and set $f=0.4$.
Lastly, to mitigate over-fitting, we incorporate the following regularization term into the loss function (6):
$\mathcal{L}_{r}(\theta)=-w_{r}\frac{1}{N_{r}}\sum^{N_{r}}_{s_{i}\sim\mathcal{S}}\zeta_{\theta}(s_{i})$ (8)
where $w_{r}$ is the a fixed regularization weight. The state $s_{i}$ is uniformly sampled from the whole state space.
2.3 Policy Learning via Constrained RL
In order to generate high-reward trajectories while satisfying the already learned constraints, an RL policy network is established and maintained during the learning process. Akin to some recent works [5, 20], this paper adopts a popular and relatively simple algorithm PPO-penalty [21] for constrained policy optimization. As shown in (9), it reshapes the original reward by turning the constraint as a penalty term into the original reward function to avoid the infeasible states, where $w_{p}$ is a fixed penalty weight and $c(s)$ is a constraint indicator defined in (1). Hence, a constrained optimal policy can be straightforwardly learned by optimizing the reshaped reward $r^{\prime}(s,a)$ with the standard deep RL algorithm PPO [21].

### v3

# Learning Constraint Network from Demonstrations via Positive-Unlabeled Learning with Memory Replay

@@ Abstract
Planning for a wide range of real-world tasks necessitates to know and write not only reward function but also constraints. While most Inverse Reinforcement Learning (IRL) works are designed to learn solely the reward, this work focuses on inferring the unknown constraints from demonstration. This paper presents a positive-unlabeled (PU) learning approach to infer a continuous, and possibly nonlinear constraint network. From a PU learning view, We treat all data in demonstrations as positive (feasible) data, and learn an optimal policy to generate high-reward-winning but potentially infeasible trajectories, which serve as unlabeled data containing both feasible and infeasible states. Under an assumption on data distribution, a feasible-infeasible classifier (i.e., constraint model) is learned from the two datasets through a postprocessing PU learning technique. Additionally, an memory replay buffer is introduced to record and reuse samples from previous iterations and prevent forgetting learned constraint. The effectiveness of the proposed method is validated in three Mujoco environments, successfully inferring continuous nonlinear constraints and outperforming the baseline method in terms of constraint accuracy and policy safety.
I Introduction
@@
To effectively plan for various robotics and automation tasks, it is essential to explicitly define the constraints that determine which states or trajectories are feasible and which should be avoided [1, 2]. Sometimes these constraints are initially unknown or hard to specify mathematically, especially when they are continuous, multivariate and inherent to an expert’s preference and experience. For example, human drivers may determine an implicit minimum distance from other cars based on traffic conditions, traffic rules, and even weather. To learn a driving policy matching human behaviors, an explicit constraint should be inferred somehow, e.g., from existing human demonstration sets.
An emerging approach to recover the underlying constraint is through Inverse Constrained Reinforcement Learning (ICRL) [3, 4, 5]. It originates from Inverse Reinforcement Learning (IRL), a method to infer a reward function from expert demonstration[6]. But in contrast to IRL, ICRL is developed exclusively to infer constraints instead of reward functions. The majority of prior works limit themselves to learning simple linear constraints, or require strong knowledge of the true constraint parameterization or environmental model. To see this, next we will review the existing constraint inferring methods. Note that given the wide range of constraint types explored in the literature, it is impractical to address them all. This work specifically focus on constraints related to avoiding certain undesirable states (or actions) throughout the trajectory (discussed in detail in section II-A). Other constraint types are out of the scope of this work.
Figure 1: The framework of the proposed method. It alternates between two steps: learning policy using Constrained RL and learning constraint from policy and demonstrations with positive-unlabeled learning.
@@
Related Works: Constraint inference has drawn more and more attention since 2018. [7] first explores how to infer constraints given an environment model and a demonstration set. They assume that all possible trajectories that could earn higher rewards than the demonstration must be constrained in some way, or the expert could have passed those trajectories. In practice, they use hit-and-run sampling to obtain such higher reward trajectories and solve an integer program to recover the constrained states in a tabular environment. [8] formalizes this idea by casting the problem in the maximum likelihood inference framework, which is also prevalent in the IRL domain. They introduce a Boltzmann policy model, where the likelihood of any feasible trajectory is assumed to be proportional to the exponential return of the trajectory, while the likelihood of any infeasible trajectory is 0. Then a greedy algorithm is proposed to add the smallest number of constraints that maximize the likelihood of the demonstrations. This framework has the advantage of being able to work with sub-optimal demonstrations. Further, [9] and [10] extend the maximum entropy framework from a deterministic setting into a stochastic setting, with soft or probabilistic constraints.
@@
Unfortunately, all the methods mentioned above only apply to systems with discrete finite state spaces. To overcome this drawback, recent papers have made a few attempts to extend this to continuous cases. [11] extends their previous mixed-integer-programming approach to learn constraint that can be expressed as (union of) axis-aligned boxes. [12] proposes an algorithm that learns a neural network policy via deep RL and generates high return trajectories with the policy network. Then the constraint is again added based on the maximum likelihood principle. [3] proposes Maximum Entropy Constraint Learning (MECL) algorithm. It not only approximates the policy with a network but also learns a constraint network, a continuous function that can represent more general constraints. This constraint network is optimized by making a gradient ascent on the maximum likelihood objective function. More recently, [4] improves MECL by proposing a variational approach that leans the distribution of constraint to capture the epistemic uncertainty. Although theoretically their maximum entropy framework can learn nonlinear constraint on continuous state spaces, in practice they was only applied to recover plane constraints, e.g. $x\geq-3$ [3], and, as we show later in the paper, performs relatively poorly at learning more nonlinear constraints. This work identifies that this limited performance results partly from two fundamental drawbacks of the maximum entropy framework: 1) the framework cannot make full use of data generated in earlier iteration and is prone to forgetting constrained region learned in earlier iterations (discussed in section II-D.), and 2) it always trains with all trajectories produced by the RL policy indiscriminately. Nevertheless, in practice some terribly learned trajectories will in fact harm the whole constraint learning.
In addition to these ICRL method, it is worth mentioning that another line of constraint inference methods have also been developed to learn continuous constraints via KKT optimality condition [13, 14]. Despite their effectiveness, these methods require a closed-form environmental model and its full derivative, which are not always available in reality. In contrast, ICRL method, including ours, only requires a simulator to generate data or can directly interact with the real world.
@@
To overcome the aforementioned challenge, especially those from the maximum entropy framework, we propose a novel positive-unlabeled approach to learn a continuous and potentially nonlinear constraint function in continuous state-action spaces. Our method is inspired by a machine learning subarea Positive-Unlabeled (PU) learning [15], and to best of our knowledge, this work is the first to handle the constraint inference problem from a PU learning perspective. Here, we treat all data from demonstrations (assumed to be feasible) as positive (i.e., feasible), and learn a policy to constantly generate potentially infeasible trajectories, which contains unlabeled states that could be feasible or infeasible. Then a PU learning technique postprocessing is applied to synthesis a feasible-infeasible classifier (i.e., constraint model) by adapting a trained labeled-unlabeled classifier with a moved classification threshold.
@@
Our learning paradigm is phrased as an iterative framework, see Fig. 1. At each iteration, we first update the policy to maximize the reward subject to the current constraints. Then, we generate a batch of trajectories from policy and only keep the high-reward-winning ones, which are deemed as potentially truly infeasible trajectories and further used along with the demonstrations to update constraints via PU learning technique. In the next iteration, a new policy is updated with respect to the new constraints.
One advantage of the proposed PU learning approach is that it can naturally utilize the data from previous iterations to learn constraint, since they can be all simply regarded as unlabeled data. To enable this, we draw inspiration from off-policy RL and introduce an memory replay buffer to record and reuse the previous data to prevent forgetting, where some most representative infeasible states in each iterations are recorded and reused for training in the following iterations. Our experiments section will prove the effectiveness of this improvement.
II Method
II-A Preliminaries and Problem Statements
@@
For a Markov decision process (MDP) [16, 17], states and actions are denoted by $s\in\mathcal{S}$ and $a\in\mathcal{A}$. $\gamma$ denotes the discount factor, and the real-valued reward is denoted by $r(s,a)$. For constraint learning with ICRL, reward is usually assumed to be known, which is usually minimum distance or minimum time [3, 7]. A trajectory $\tau=\{s_{1},a_{1},\ldots,s_{T},a_{T}\}$ contains a sequence of state-action pairs in one episode. For the trajectory $\tau$, the total discounted reward (return) is defined as $r(\tau)=\sum_{t=1}^{T}\gamma^{t}r(s_{t},a_{t})$. A policy, the mapping between states and actions, is denoted by $\pi(a|s)$. Note that different from [14, 13], we do not make the assumption of possessing a closed-form model or its derivatives of the transition dynamics. We only require a simulator that can simulate the dynamics and return the next state and reward.
@@
The true constraint set $\mathcal{C^{*}}$ is defined as the set of all actually infeasible states $\mathcal{C^{*}}=\{s\in\mathcal{S}|s\texttt{ is actually infeasible}\}$. Note that for simplicity of explanation, we only write the state constraint, but it can be easily extended to state-action constraints by augmenting the state with the action to form an augmented state. We aim to recover the true constraint set from the demonstration. Suppose we have collected a set of demonstrated trajectories $\mathcal{D}=\{\tau^{i}\}^{M_{d}}_{i=1}$ generated from an expert $\pi^{*}$ navigating in the true environment. We assume that the expert $\pi^{*}$ maximizes the return $J(\pi)=\mathbb{E}_{\tau\sim\pi}\{r(\tau)\}$ while never visiting the truly infeasible states. To represent and learn an arbitrary constraint set in continuous state space, we define a continuous constraint function $\zeta_{\theta}(s)\in(0,1)$ and its induced constraint set $\mathcal{C}_{\theta}=\{s\in\mathcal{S}|\zeta_{\theta}(s)\leq d\}$, where $\theta$ is the function parameter and $d$ is the classification threshold. The constraint function $\zeta_{\theta}$ can be regarded as the probability that a state is feasible.
II-B Constraint Inference as Positive-Unlabeled Learning
@@
We extract the underlying constraint by contrasting the feasible demonstrations with some potentially infeasible trajectories generated by a reward-maximizing policy. In each iteration, we first sample a set of trajectories $\mathcal{P}=\{\tau_{p}^{i}\}^{M_{p}}_{i=1}$ by performing current policy $\pi_{\phi}$ (discussed later in section II-C). Each sampled trajectory consists of sequential state-action pairs $\tau_{p}=\{s_{j},a_{j}\}_{j=1}^{N_{i}}$. The trajectories winning higher reward than demonstrator is considered to having violating some unknown constraints. However, it remains unclear which specific state(s) within the trajectory has violated the constraint. In another word, trajectory $\tau_{p}$ consist of both feasible states and infeasible states but both remains unlabeled. In contrast, it is certain that all states on the demonstrated trajectories are labeled as feasible. Our goal is to classify each single state as feasible or infeasible by learning from a batch of fully labeled feasible states and another batch of unlabeled states.
This insight inspires us to formulate constraint inference as a positive-unlabeled learning problem, which learns from a positive dataset, and an unlabeled dataset containing both positive and negative samples [15]. Within our framework, demonstrations and feasible states from them are designated as positive, while the policy offers unlabeled samples with unknown feasibility. Any datapoint can be represented as a set of triplets $(s,c,l)$ with $s$ the state, $c$ the true class (1 for feasible), and $l$ a binary variable representing whether the tuple is labeled. If a sample is labeled (i.e., $l=1$), it must originate from the demonstration and is sure to be feasible, i.e., $Pr(c=1|l=1)=1$. A key quantity from PU learning is the label frequency $f$, which is defined as the proportion of positive samples that are labeled within the entire dataset:
@@
In the constraint learning context, the label frequency $f$ implies the density of the truly feasible states in the state space. A high $f$ suggests that the majority of truly feasible states are distributed solely on the demonstrated trajectories and thus labeled, and the undemonstrated state regions are mostly infeasible. Subsequently, we introduce an assumption regarding the labeling mechanism [15]:
@@
Given the SCAR assumption, we introduce a postprocessing PU learning method [18] for uncovering the unknown infeasible regions due to two advantages: 1) inherent simplicity and 2) compatibility with various models, including neural networks. Based on SCAR assumption, the probability of an sample being labeled is directly proportional to the probability of that sample being positive:
@@
Consequently, it is evident that a labeled-unlabeled classifier $Pr(l|s)$, trained by treating unlabeled data as negative and labeled data as positive, can be directly employed to forecast the feasibility probabilities $Pr(c=1|s)$ and thus determine its feasibility [18]. Alternatively, noticing that $Pr(c=1|s)>0.5$ is equivalent to $Pr(l=1|s)>0.5f$, the labeled-unlabeled classifier can be repurposed as a feasible-infeasible classifier by directly adjusting the decision threshold from $d=0.5$ to $d=0.5f$.
@@
Once trained, this classifier can be directly employed as a feasible-infeasible classifier by setting the classification threshold to $d=0.5f$ as in (6).
@@
The subsequent work involves determining the label frequency $f$. Estimating $f$ is a central concern in PU learning, with diverse methods such as partial matching and kernel embedding being proposed [15]. In this work, we suggest two approaches. A conventional and straightforward approach [19] uses the insight that $Pr(l=1|s)=fPr(c=1|s)$, which is equal to the label frequency $f$ when the true class probability is $Pr(c=1|s)=1$. Suppose there will be states for which $Pr(c=1|s)\approx 1$, the label frequency can be hence estimated as $f=\max_{s\in\mathcal{D}}\zeta_{\theta}(s)$. Alternatively, recall that $f$ indicates the density of infeasible states in the state space, it can be reviewed as a hyper-parameter representing the user’s belief over the constraint density. A higher value of $f$ suggests a denser distribution of infeasible states. In our experiments, we treat $f$ as a tunable parameter and report the corresponding threshold $d$ in the appendix.
II-C Policy Learning via Constrained RL
In order to generate high-reward trajectories while satisfying the already learned constraints, an RL policy network is established and maintained during the learning process. Akin to some recent works [5, 20], this paper adopts a relatively simple algorithm PPO-penalty [21] for constrained policy optimization. As shown in (7), it reshapes the original reward by turning the constraint as a penalty term into the original reward function to avoid the infeasible states, where $w_{p}$ is a fixed penalty weight and $c(s)$ is a constraint indicator defined in (6).